I. Find the present and future value of $1000 received every month end for 20 years if the interest rate is J12 = 12% p.a.
1
Expert's answer
2020-06-01T20:07:42-0400

Present value of the annuity is:

"PV = 1000*\frac{1 - (1 + 0.12)^{-20}}{0.12} = 7,469.44."

The future value of the annuity is:

"FV = 1000*\frac{(1 + 0.12)^{20} - 1}{0.12} = 72,052.44."